if i give you two points on a graph say (5,20) and (8,29) where do i go if i do
20-29=-9
--------
5-8=-3 then i do
-9 / -3 getting 3 ok then i make a table knowing that the slope is 3? but then 0 or y intercept is 5 right but then they want me to solve algebra method where do i go next from there? do i make a graph or a rule first? because the rule would be y=5x-3 right?

Answer by Alan3354(30993)

if i give you two points on a graph say (5,20) and (8,29) where do i go if i do
20-29=-9
--------
5-8=-3 then i do
-9 / -3 getting 3 ok then i make a table knowing that the slope is 3? but then 0 or y intercept is 5 right but then they want me to solve algebra method where do i go next from there? do i make a graph or a rule first? because the rule would be y=5x-3 right?
-----------------
You got the slope, m, right, m = 3.
Then, use either point to get the equation.
y-y1 = m*(x-x1) (x1,y1) is the point.
y-20 = 3*(x-5)
y-20 = 3x-15
y = 3x + 5
The y-intercept is 5, at (0,5).
--------------------------
Then you can make one more point, and since it's a straight line, draw the line thru them.
Give x any value, and find y.
eg, x = 3, y = 14 --> (3,14)